# DDAD: A Two-pronged Adversarial Defense Based on Distributional Discrepancy

### v1

@@ 4.2 Training Process of DDAD
$:=\widehat{\textsc{MMD}}^{2}_{\rm u}(S_{\mathcal{C}},S_{\mathcal{A}};k_{\omega})/\hat{\sigma}_{\lambda}(S_{\mathcal{C}},S_{\mathcal{A}};k_{\omega}),$ (5)
where $S_{\mathcal{C}}$ are clean samples, $S_{\mathcal{A}}$ can be any adversarial samples, $\hat{\sigma}^{2}_{\lambda}$ is a regularized estimator of $\sigma^{2}$ and $\lambda$ is a small constant to avoid 0 division (here we assume $m=n$ to obtain the asymptotic distribution of the MMD estimator):
$\hat{\sigma}^{2}_{\lambda}$ $=\frac{4}{n^{3}}\sum^{n}_{i=1}\left(\sum^{n}_{j=1}H_{ij}\right)^{2}$
$-\frac{4}{n^{4}}\left(\sum^{n}_{i=1}\sum^{n}_{j=1}H_{ij}\right)^{2}+\lambda.$ (6)
We can obtain optimized $k_{\omega}$ (we denote it as $k^{*}_{\omega}$) by maximizing Eq. 4.2 on the training set. Then, we define MMD-OPT as the MMD estimator with an optimized characteristic kernel $k^{*}_{\omega}$:
@@
Training the denoiser. In this paper, we use DUNET (Liao et al., 2018) as our denosing model. To train the denoiser, we first randomly generate noise $\mathbf{n}$ from a Gaussian distribution $\mathbb{N}(\mu,\sigma^{2})$ and add $\mathbf{n}$ to $S_{\mathcal{A}}$ that are generated from clean data-label pairs $(S_{\mathcal{C}},Y_{\mathcal{C}})$, resulting in noise-injected AEs:
@@
4.3 Inference Process of DDAD
In this section, we discuss the two-pronged inference process of DDAD in detail.
The use of validation data. In the inference phase, we define a batch of clean validation data as $S_{\mathcal{V}}$ and the test data as $S_{\mathcal{T}}$. In practice, $S_{\mathcal{V}}$ is extracted from the training data and is completely inaccessible during the training. Then $S_{\mathcal{V}}$ serves as a reference to measure the distributional discrepancy. According to Eq. 7, the distributional discrepancies between $S_{\mathcal{V}}$ and $S_{\mathcal{T}}$ can be defined as:
@@ 4.3 Inference Process of DDAD
The two-pronged inference process. (1) if MMD-OPT($S_{\mathcal{V}}$, $S_{\mathcal{T}}$) in Eq. 9 is less than a threshold value $t$, i.e., MMD-OPT($S_{\mathcal{V}}$, $S_{\mathcal{T}}$) $<t$, then $S_{\mathcal{T}}$ will be treated as CEs and directly fed into the classifier. Then the output will be $\widehat{h^{*}_{\mathcal{C}}}(S_{\mathcal{T}})$, where $\widehat{h^{*}_{\mathcal{C}}}$ is a well-trained classifier; (2) otherwise, $S_{\mathcal{T}}$ will be treated as AEs and denoised by the denoiser. Then, the output will be $\widehat{h^{*}_{\mathcal{C}}}(g_{\bm{\theta}^{*}}(S_{\mathcal{T}}))$, where $g_{\bm{\theta}^{*}}$ is a well-trained denoiser by our method.
@@
Ours DDAD 78.61 $\pm$ 0.04 53.85 $\pm$ 0.23
5 Experiments
5.1 Experiment Settings
We briefly introduce the experiment settings here and provide a more detailed version in Appendix D.
Dataset and target models. We evaluate DDAD on two benchmark datasets with different scales, i.e., CIFAR-10 (Krizhevsky et al., 2009) and ImageNet-1K (Deng et al., 2009). For the target models, we mainly use three architectures with different capacities: ResNet (He et al., 2016), WideResNet (Zagoruyko & Komodakis, 2016) and Swin Transformer (Liu et al., 2021).
Baseline settings. DDAD is a two-pronged adversarial defense method, which is different from most existing defense methods. In terms of the pipeline structure, MagNet (Meng & Chen, 2017) is the only similar defense method to ours, which also contains a two-pronged process. However, MagNet is now considered outdated, making it unfair for DDAD to compare with it. Therefore, to make the comparison as fair as possible, we follow a recent study on robust evaluation (Lee & Kim, 2023) to compare our method with SOTA adversarial training (AT) methods in RobustBench (Croce et al., 2020) and adversarial purification (AP) methods selected by Lee & Kim (2023).
Evaluation settings. We mainly use PGD+EOT (Athalye et al., 2018b) and AutoAttack (Croce & Hein, 2020a) to compare our method with different baseline methods. Specifically, following Lee & Kim (2023), we evaluate AP methods on the PGD+EOT attack with 200 PGD iterations for CIFAR-10 and 20 PGD iterations for ImageNet-1K. We set the EOT iteration to 20 for both datasets. We evaluate AT baseline methods using AutoAttack with 100 update iterations, as AT methods have seen PGD attacks during training, leading to overestimated results when evaluated on PGD+EOT (Lee & Kim, 2023). For our method, we implicitly design an adaptive white-box attack by considering the entire defense mechanism of DDAD (see Algorithm 3). To make a fair comparison, we evaluate our method on both adaptive white-box PGD+EOT attack and adaptive white-box AutoAttack with the same configurations mentioned above. Notably, we find that our method achieves the worst-case robust accuracy on adaptive white-box PGD+EOT attack. Therefore, we report the robust accuracy of our method on adaptive white-box PGD+EOT attack for Table 1 and 2. On CIFAR-10, the maximum allowed perturbaiton budget $\epsilon$ for $\ell_{\infty}$-norm-based attacks and $\ell_{2}$-norm-based attacks is set to $8/255$ and $0.5$, respectively. While on ImageNet-1K, we set $\epsilon=4/255$ for $\ell_{\infty}$-norm-based attacks.
Implementation details of DDAD. To avoid the evaluation bias caused by seeing similar attacks beforehand during training, we train both the MMD-OPT and the denoiser using $\ell_{\infty}$-norm MMA attack (Gao et al., 2022), which differs significantly from PGD+EOT and AutoAttack. Then, we use unseen attacks to evaluate DDAD. For optimizing the MMD, following Gao et al. (2021), we set the learning rate to be $2\times 10^{-4}$ and the epoch number to be 200. For training the denoiser, we set the epoch number to be 60. The initial learning rate is set to $1\times 10^{-3}$ for both datasets and is divided by 10 at the 45th and 60th epoch to avoid robust overfitting (Rice et al., 2020). We provide more detailed implementation details in Appendix D.
@@ 5.2 Defending against Adaptive White-box Attacks
Result analysis on CIFAR-10. Table 1 shows the evaluation performance of DDAD against adaptive white-box PGD+EOT attack with $\ell_{\infty}(\epsilon=8/255)$ and $\ell_{2}(\epsilon=0.5)$ on CIFAR-10. Compared to SOTA defense methods, DDAD improves clean and robust accuracy by a notable margin. The evaluation results against BPDA+EOT on CIFAR-10 can be found in Appendix E.1.
Result analysis on ImageNet-1K. Table 2 shows the evaluation performance of DDAD against adaptive white-box PGD+EOT attack with $\ell_{\infty}(\epsilon=4/255)$ on ImageNet-1K. The advantages of our method over baselines become more significant on large-scale datasets. Specifically, compared with AP methods that rely on density estimation (Nie et al., 2022; Lee & Kim, 2023), our method improves clean accuracy by at least 7.13% and robust accuracy by 11.70% on ResNet-50. This empirical evidence supports that identifying distributional discrepancies is a simpler and more feasible task than estimating data density, especially on large-scale datasets such as ImageNet-1K.
@@ 5.3 Defending against Unseen Transfer Attacks
Since DDAD requires AEs to train the MMD-OPT and the denoiser, it is important for us to evaluate the transferability of our method. We report the transferability of our method (trained on WideResNet-28-10) under different threat models, which include WideResNet-70-16, ResNet-18, ResNet-50 and Swin Transformer in Appendix E.2. We use PGD+EOT $\ell_{\infty}$ and C&W $\ell_{2}$ (Carlini & Wagner, 2017) with 200 iterations for evaluation. Experiment results show that our method generalizes well to unseen transfer attacks.
@@ 5.4 Ablation Studies
Ablation study on mixed data batches. We explore a more challenging scenario for our method, in which each data batch contains a mixture of CEs and AEs. Figure 2 (bottom) shows the mixed accuracy (i.e., the accuracy on mixed data) of our method on CIFAR-10 with different proportions of AEs (generated by adaptive white-box PGD+EOT $\ell_{\infty}$ with $\epsilon=8/255$) in each batch, ranging from 0% (i.e., pure CEs) to 100% (i.e., pure AEs). Initially, the mixed accuracy drops from over 90% to approximately 80%. This is because, with a high proportion of CEs, the MMD-OPT has a high chance to regard the entire batch as clean data. After that (i.e., from 30% onwards), the mixed accuracy degrades gradually to approximately 70%. This is because, as the proportion of AEs increases, the MMD-OPT regards the entire batch as adversarial and feeds it into the denoiser. Notably, DDAD can still outperform baseline methods (see Appendix E.3).
Ablation study on threshold values of MMD-OPT. We explore the impact of threshold values of MMD-OPT in Appendix E.4. We select the threshold based on the experimental results on the validation data. Specifically, a threshold value of $0.5$ is selected for CIFAR-10 and $0.02$ is selected for ImageNet-1K. It is reasonable to use a smaller threshold for ImageNet-1K because the distribution of AEs with $\epsilon=4/255$ (i.e., AEs for ImageNet-1K) will be closer to CEs than AEs with $\epsilon=8/255$ (i.e., AEs for CIFAR-10). Intuitively, when $\epsilon$ decreases to 0, AEs are the same as CEs (i.e., the distribution of AEs and CEs will be the same).
Ablation study on injecting Gaussian noise. We provide evaluation results of our method against adaptive white-box PGD+EOT attack with and without injecting Gaussian noise on CIFAR-10 in Appendix E.5. We find that injecting Gaussian noise can make DDAD generalize better.
Ablation study on the two-pronged process. We provide evaluation results of our method against adaptive white-box PGD+EOT attack with and without MMD-OPT on CIFAR-10 in Appendix E.6. We find that using the two-pronged process can largely improve clean accuracy.
@@ 5.5 Combination with Adversarial Training
We also combine several well-known AT methods (Madry et al., 2018; Zhang et al., 2019; Wang et al., 2020; Wu et al., 2020) with DDAD: detected CEs are directly classified by the AT-based classifier, while detected AEs are first denoised by our method before being classified. Experimental results show that combining AT methods with DDAD further enhances adversarial robustness (see Appendix E.7).
5.6 Compute Resource of DDAD
We report the compute resources used for training and evaluating DDAD in Appendix E.8. Compared to AT baselines, DDAD offers better training efficiency (e.g., it can scale to large datasets like ImageNet-1K). Additionally, although DDAD requires training an extra denoiser and MMD-OPT, it significantly outperforms AP baselines in inference speed. Furthermore, relying on a pre-trained generative model is not always feasible, as training such models at scale can be highly resource-intensive. Therefore, in general, DDAD provides a more lightweight design.

### v2

@@ 4.2 Training Process of DAD
$:=\widehat{\textsc{MMD}}^{2}_{\rm u}(S_{\mathcal{C}},S_{\mathcal{A}};k_{\omega})/\hat{\sigma}_{\lambda}(S_{\mathcal{C}},S_{\mathcal{A}};k_{\omega}).$ (4)
Here $S_{\mathcal{C}}$ are clean samples, $S_{\mathcal{A}}$ are adversarial samples, $\hat{\sigma}^{2}_{\lambda}$ is a regularized estimator of $\sigma^{2}$:
$\frac{4}{n^{3}}\sum^{n}_{i=1}\left(\sum^{n}_{j=1}H_{ij}\right)^{2}-\frac{4}{n^{4}}\left(\sum^{n}_{i=1}\sum^{n}_{j=1}H_{ij}\right)^{2}+\lambda,$ (5)
where $\lambda$ is a small constant to avoid 0 division (here we assume $m=n$ to obtain the asymptotic distribution of the MMD estimator).
We can obtain optimized $k_{\omega}$ (we denote it as $k^{*}_{\omega}$) by maximizing Eq. (4.2) on the training set. Then, we define MMD-OPT as the MMD estimator with the optimized kernel $k^{*}_{\omega}$:
@@
3:  for epoch $=1,...,T$ do
5:    $S^{\prime}_{\mathcal{A}}\leftarrow$ AEs generated from $(S^{\prime}_{\mathcal{C}},Y^{\prime}_{\mathcal{C}})$;
6:    generate Gaussian noise: $\mathbf{n}\sim\mathbb{N}(\mu,\sigma^{2})$;
7:    $S^{\prime}_{\text{noise}}\leftarrow S^{\prime}_{\mathcal{A}}+\mathbf{n}$;
[…]
9:    Update $\bm{\theta}$ by Eq. (4.2);
10:  end for
11:  Output: denoiser $g$ with well-trained parameters $\bm{\theta}^{*}$
Algorithm 2 Training the denoiser.
First bird: MMD-OPT-based denoiser. In this paper, we use DUNET (Liao et al., 2018) as our denoising model. To train the denoiser, we first randomly generate noise $\mathbf{n}$ from a Gaussian distribution $\mathbb{N}(\mu,\sigma^{2})$ and add $\mathbf{n}$ to $S_{\mathcal{A}}$ that are generated from clean data-label pairs $(S_{\mathcal{C}},Y_{\mathcal{C}})$, resulting in noise-injected AEs:
@@
4.3 Inference Process of DAD
In this section, we demonstrate the two-pronged inference process of DAD in detail.
Second bird: discriminator based on MMD-OPT. In the inference phase, we define a batch of clean validation data as $S_{\mathcal{V}}$ and the test data as $S_{\mathcal{T}}$. In practice, $S_{\mathcal{V}}$ is extracted from the training data and is completely inaccessible during the training. Then $S_{\mathcal{V}}$ serves as a reference to measure the distributional discrepancy. According to Eq. (6), the distributional discrepancies between $S_{\mathcal{V}}$ and $S_{\mathcal{T}}$ can be
@@ 4.3 Inference Process of DAD
[…]
The two-pronged inference process. Based on the denoiser and the discriminator based on MMD-OPT, the two-pronged inference process is demonstrated below.
  • If MMD-OPT($S_{\mathcal{V}}$, $S_{\mathcal{T}}$) in Eq. (8) is less than a threshold $t$, i.e., MMD-OPT($S_{\mathcal{V}}$, $S_{\mathcal{T}}$) $<t$, then $S_{\mathcal{T}}$ will be treated as CEs and directly fed into the classifier. Then the output will be $\widehat{h^{*}_{\mathcal{C}}}(S_{\mathcal{T}})$, where $\widehat{h^{*}_{\mathcal{C}}}$ is a well-trained classifier;
  • Otherwise, $S_{\mathcal{T}}$ will be treated as AEs and denoised by the well-trained denoiser $g_{\bm{\theta}^{*}}$. Then, the output will be $\widehat{h^{*}_{\mathcal{C}}}(g_{\bm{\theta}^{*}}(S_{\mathcal{T}}))$.
5 Experiments
We briefly introduce the experiment settings here and provide a more detailed version in Appendix C.
Dataset and target models. We evaluate DAD on two benchmark datasets with different scales, i.e., CIFAR-10 (Krizhevsky et al., 2009) and ImageNet-1K (Deng et al., 2009). For the target models, we mainly use three architectures with different capacities: ResNet (He et al., 2016), WideResNet (Zagoruyko & Komodakis, 2016) and Swin Transformer (Liu et al., 2021b).
Baseline settings. DAD is a two-pronged adversarial defense method, which is different from most existing defense methods. In terms of the pipeline structure, MagNet (Meng & Chen, 2017) is the only similar defense method to ours, which also contains a two-pronged process. However, MagNet is now considered outdated, making it unfair for DAD to compare with it. Therefore, to make the comparison as fair as possible, we follow a recent study on robust evaluation (Lee & Kim, 2023) to compare our method with SOTA adversarial training (AT) methods in RobustBench (Croce et al., 2020) and adversarial purification (AP) methods selected by Lee & Kim (2023).
Algorithm 3 Adaptive white-box PGD+EOT attack for DAD.
3:  for PGD iteration $1,...,T$ do
[…]
6:    for EOT iteration $1,...,K$ do
[…]
9:      else
10:         Generate Gaussian noise: $\mathbf{n}\sim\mathbb{N}(\mu,\sigma^{2})$;
11:         $S_{\text{noise}}\leftarrow S_{\mathcal{A}}+\mathbf{n}$;
12:         ${\mathcal{G}}_{\text{EOT}}\leftarrow{\mathcal{G}}_{\text{EOT}}+\nabla_{S_{\mathcal{A}}}(\text{MMD-OPT}(S_{\mathcal{C}},S_{\mathcal{A}})+\alpha\cdot{\mathcal{L}}_{\text{ce}}(\widehat{h_{\mathcal{C}}^{*}}(g_{\bm{\theta}}(S_{\text{noise}})),Y_{\mathcal{C}}))$;
13:      end if
14:    end for
15:    ${\mathcal{G}}_{\text{EOT}}\leftarrow\frac{1}{K}{\mathcal{G}}_{\text{EOT}}$;
16:    Update $S_{\mathcal{A}}\leftarrow\Pi_{\mathcal{B}_{\epsilon}(S_{\mathcal{C}})}\left(S_{\mathcal{A}}+\eta\cdot\text{sign}({\mathcal{G}}_{\text{EOT}})\right)$;
17:  end for
18:  Output: $S_{\mathcal{A}}$
Evaluation settings. We mainly use PGD+EOT (Athalye et al., 2018b) and AutoAttack (Croce & Hein, 2020a) to compare our method with different baseline methods. Specifically, following Lee & Kim (2023), we evaluate AP methods on the PGD+EOT attack with 200 PGD iterations for CIFAR-10 and 20 PGD iterations for ImageNet-1K. We set the EOT iteration to 20 for both datasets. We evaluate AT baseline methods using AutoAttack with 100 update iterations, as AT methods have seen PGD attacks during training, leading to overestimated results when evaluated on PGD+EOT (Lee & Kim, 2023). For our method, we implicitly design an adaptive white-box attack by considering the entire defense mechanism of DAD (see Algorithm 3). To make a fair comparison, we evaluate our method on both adaptive white-box PGD+EOT attack and adaptive white-box AutoAttack with the same configurations mentioned above. Notably, we find that our method achieves the worst-case robust accuracy on adaptive white-box PGD+EOT attack. Therefore, we report the robust accuracy of our method on adaptive white-box PGD+EOT attack for Table 1 and 2. On CIFAR-10, the maximum allowed perturbation budget $\epsilon$ for $\ell_{\infty}$-norm-based attacks and $\ell_{2}$-norm-based attacks is set to $8/255$ and $0.5$, respectively. While on ImageNet-1K, we set $\epsilon=4/255$ for $\ell_{\infty}$-norm-based attacks.
Implementation details of DAD. To avoid the evaluation bias caused by seeing similar attacks beforehand during training, we train both the MMD-OPT and the denoiser using $\ell_{\infty}$-norm MMA attack (Gao et al., 2022), which differs significantly from PGD+EOT and AutoAttack. Then, we use unseen attacks to evaluate DAD. For optimizing the MMD, following Gao et al. (2021), we set the learning rate to be $2\times 10^{-4}$ and the epoch number to be 200. For training the denoiser, we set the epoch number to be 60. The initial learning rate is set to $1\times 10^{-3}$ for both datasets and is divided by 10 at the 45th and 60th epoch to avoid robust overfitting (Rice et al., 2020). We provide more detailed implementation details in Appendix C.
@@ 5 Experiments
Ours DAD 78.61 $\pm$ 0.04 53.85 $\pm$ 0.23
@@ 5.1 Defending against Adaptive White-box Attacks
Result analysis on CIFAR-10. Table 1 shows the evaluation performance of DAD against adaptive white-box PGD+EOT attack with $\ell_{\infty}(\epsilon=8/255)$ and $\ell_{2}(\epsilon=0.5)$ on CIFAR-10. Compared to SOTA defense methods, DAD improves clean and robust accuracy by a notable margin. The evaluation results against BPDA+EOT on CIFAR-10 can be found in Appendix D.1.
Result analysis on ImageNet-1K. Table 2 shows the evaluation performance of DAD against adaptive white-box PGD+EOT attack with $\ell_{\infty}(\epsilon=4/255)$ on ImageNet-1K. The advantages of our method over baselines become more significant on large-scale datasets. Specifically, compared with AP methods that rely on density estimation (Nie et al., 2022; Lee & Kim, 2023), our method improves clean accuracy by at least 7.13% and robust accuracy by 11.70% on ResNet-50. This empirical evidence supports that identifying distributional discrepancies is a simpler and more feasible task than estimating data density, especially on large-scale datasets such as ImageNet-1K.
@@ 5.2 Defending against Unseen Transfer Attacks
Since DAD requires AEs to train the MMD-OPT and the denoiser, it is important for us to evaluate the transferability of our method. We report the transferability of our method (trained on WideResNet-28-10) under different threat models, which include WideResNet-70-16, ResNet-18, ResNet-50 and Swin Transformer in Table 3. We use PGD+EOT $\ell_{\infty}$ and C&W $\ell_{2}$ (Carlini & Wagner, 2017) with 200 iterations for evaluation. Experiment results show that our method generalizes well to unseen transfer attacks.
Table 3: Robust accuracy (%) of DAD trained on WideResNet-28-10 against unseen transfer attacks on CIFAR-10. Attackers cannot access parameters of WideResNet-28-10, thereby it is in a gray-box setting. We report averaged results and standard deviations of 5 runs.
Trained on WRN-28-10
Unseen Transfer Attack WRN-70-16 RN-18 RN-50 Swin-T
@@ 5.3 Ablation Studies
Ablation study on mixed data batches. We explore a more challenging scenario for our method, in which each data batch contains a mixture of CEs and AEs. Figure 2 (bottom) shows the mixed accuracy (i.e., the accuracy on mixed data) of our method on CIFAR-10 with different proportions of AEs (generated by adaptive white-box PGD+EOT $\ell_{\infty}$ with $\epsilon=8/255$) in each batch, ranging from 0% (i.e., pure CEs) to 100% (i.e., pure AEs). Initially, the mixed accuracy drops from over 90% to approximately 80%. This is because, with a high proportion of CEs, the MMD-OPT has a high chance to regard the entire batch as clean data. After that (i.e., from 30% onward), the mixed accuracy degrades gradually to approximately 70%. This is because, as the proportion of AEs increases, the MMD-OPT regards the entire batch as adversarial and feeds it into the denoiser. Notably, DAD can still outperform baseline methods (see Appendix D.2).
Ablation study on threshold values of MMD-OPT. We explore the impact of threshold values of MMD-OPT in Appendix D.3. We select the threshold based on the experimental results on the validation data. Specifically, a threshold value of $0.5$ is selected for CIFAR-10 and $0.02$ is selected for ImageNet-1K. It is reasonable to use a smaller threshold for ImageNet-1K because the distribution of AEs with $\epsilon=4/255$ (i.e., AEs for ImageNet-1K) will be closer to CEs than AEs with $\epsilon=8/255$ (i.e., AEs for CIFAR-10). Intuitively, when $\epsilon$ decreases to 0, AEs are the same as CEs (i.e., the distribution of AEs and CEs will be the same).
Ablation study on the regularization term $\alpha$ in Eq. (4.2). We explore the impact of $\alpha$ in Appendix D.4. Notably, when $\alpha$ increases, the robust accuracy will decrease. This is because increasing $\alpha$ in Eq. (4.2) reduces the relative influence of the MMD-OPT, thereby diminishing its contribution to robustness. This observation aligns with the theoretical findings presented in Section 3.
Ablation study on injecting Gaussian noise. We provide evaluation results of our method against adaptive white-box PGD+EOT attack with and without injecting Gaussian noise on CIFAR-10 in Appendix D.5. We find that injecting Gaussian noise can make DAD generalize better.
Ablation study on the two-pronged process. We provide evaluation results of our method against adaptive white-box PGD+EOT attack with and without MMD-OPT on CIFAR-10 in Appendix D.6. We find that using the two-pronged process can largely improve clean accuracy.
5.4 Compute Resource of DAD
We report the compute resources used for training and evaluating DAD in Appendix D.7. Compared to AT baselines, DAD offers better training efficiency (e.g., it can scale to large datasets like ImageNet-1K). Additionally, although DAD requires training an extra denoiser and MMD-OPT, it significantly outperforms AP baselines in inference speed. Furthermore, relying on a pre-trained generative model is not always feasible, as training such models at scale can be highly resource-intensive. Therefore, in general, DAD provides a more lightweight design.
@@ 5.5 Combination with Adversarial Training
We also combine several well-known AT methods with DAD: Vanilla AT (Madry et al., 2018), TRADES (Zhang et al., 2019), MART (Wang et al., 2020), and MART (Wu et al., 2020) to see whether our method can be combined with AT-based methods. Following our pipeline, detected CEs are directly classified by the AT-based classifier, while detected AEs are first denoised by our method before being classified by AT-based methods. Experimental results show that combining AT methods with DAD can consistently improve adversarial robustness under different proportions of AEs in each batch (see Appendix D.8).